The train 75 km apart approach each other en parallel tracks, each moving at
kmh. A bird flies back and forth between the trains at 20km/h until the trains pass
each other. How for does the bird fly
 
 
 

Grade:12th pass

1 Answers

Unicorn
2 years ago
Since You have not mentioned velocity of train .assuming ot to be 75 km/hr 
Distance between trains is 75 km so time taken by trains to cross each other will be distance /relative speed =75/150 =0.5 hour 
 
Speed of bird is 20 km /hr .so distance bird can travel =speed *time =20*0.5 =10 km